Circular No. 6619

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 B. Soifer, G. Neugebauer, L. Armus, M. Metzger, S. Kulkarni, S. Djor-

govski, and C. Steidel, California Institute of Technology; and D. Frail, National Radio Astronomy Observatory, communicate: “Near-infrared images of the region near the optical variable reported by Groot et al. (IAUC 6584) were obtained on Mar. 30 and 31 UT with the Keck I 10-m telescope (+ NIRC). A source is detected in both bands J and K within 0′′.3 of the extended source reported by Groot et al. and Metzger et al. (IAUC 6588). Within a 1′′.2 aperture, the source has J = 23.5 ± 0.2 (combined data from Mar. 30.3 and 31.2), and K = 22.0 ± 0.2 (Mar. 30.2). This corresponds to a region including the point source detected by HST (IAUC 6606). A comparison with the Mar. 17.8 J magnitude reported on IAUC 6611 indicates that the near-infrared brightness of the region has strongly varied over the interval.”

K. Sahu, M. Livio, L. Petro, and F. D. Macchetto, Space Telescope Science Institute; J. van Paradijs, University of Amsterdam; and C. Kouveliotou, G. Fishman, and and C. Meegan, Marshall Space Flight Center, NASA, communicate: “GRB 970228 was observed again with the HST Wide Field and Planetary Camera (WFPC2) between Apr. 7.15 and 7.32 UT, ∼ 38 days after the outburst. The optical transient (IAUC 6584, 6588, 6606) was placed at the center of the Planetary Camera field-of-view. Observations were taken in the F606W (wide V) and F814W (I) filters. Four exposures with a total integration time of 4700 s were taken in the F606W filter and two exposures with a total integration time of 2400 s were taken in the F814W filter. The optical counterpart, as well as the extended source, were clearly detected in both passbands. The derived magnitudes for the point source are 26.0 ± 0.3 and 24.6 ± 0.3 in the F606W and F814W filters, respectively. The data were put in the HST archive on the day of the observations. The data can be accessed through the World Wide Web at

http://archive.stsci.edu/.”

TIME ADJUSTMENT ON 1997 JUNE 30 Bulletin C13 of the International Earth Rotation Service announces

that a positive leap second will be introduced such that the sequence of UTC second markers will be: 1997 June 30^{d}23^{h}59^{m}59^{s}, 30^{d}23^{h}59^{m}60^{s}, July 1^{d}00^{h}00^{m}00^{s}. Beginning 1997 July 1, the difference UTC–TAI = −31^{s}.

1997 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green